Find the probanility of getting.

a.fewer than 4 head.

b.exactly 4 head.

C.not more than 4 head.

d.two or more heads.


1
Expert's answer
2022-06-13T15:21:42-0400

Solution:

Let we flip coin 4 times. First we need to find probability of 0-head, 1-head, 2-head, 3-head and 4-head.

Number of possible outcomes: "N=2^{n};" here "n" - number of flips.

Number of ways: "^{n}C_{m}=\frac{n!}{(n-m)!m!};" here "m" - number of heads. So,

"^{4}C_{0}=\frac{4!}{(4-0)!0!}=1;"

"^{4}C_{1}=\frac{4!}{(4-1)!1!}=4;"

"^{4}C_{2}=\frac{4!}{(4-2)!2!}=6;"

"^{4}C_{3}=\frac{4!}{(4-3)!3!}=4;"

"^{4}C_{4}=\frac{4!}{(4-4)!4!}=1;"

"N=2^{4}=16;"

"P(0)=\frac{1}{16}=0.0625" - probability of 0 head;

"P(1)=\frac{4}{16}=0.25" - probability of 1 head;

"P(2)=\frac{6}{16}=0.375" - probability of 2 heads;

"P(3)=\frac{4}{16}=0.25" - probability of 3 heads;

"P(4)=\frac{1}{16}=0.0625" - probability of 4 heads.

a) Fewer than 4 heads:

"P(m<4)=P(0)+P(1)+P(2)+P(3)=0.0625+0.25+0.375+0.25=0.9375;"

b) Exactly 4 heads:

"P(m=4)=P(4)=0.0625;"

c) Not more than 4 heads:

"P(m<=4)=P(0)+P(1)+P(2)+P(3)+P(4)=0.0625+0.25+0.375+0.25+0.0625=1;"

d) Two or more heads:

"P(m>=2)= P(2)+P(3)+P(4)=0.375+0.25+0.0625=0.6875."

Answer:

a) "P(m<4)=93.75\%;"

b) "P(m=4)=6.25\%;"

c) "P(m<=4)=100\%;"

d) "P(m>=2)=68.75\%;"